The time taken by a particle executing SHM of time period T to move from the mean position to half the maximum displacement is,

Grade:11

2 Answers

Sanjeev Malik IIT BHU
11 years ago

let eqn be x= Asin(wt)

now A/2=Asin(wt)  wherw w=2pie/T

 

solve to get t=T/12

 

please approve and like the answer by clicking "yes" below

Kushagra Madhukar
3 years ago
Dear student,
Please find the answer to your problem below.
 
let eqn of SHM be x = Asin(wt)
now A/2=Asin(wt)  (where w=2pi/T)
solve to get t=T/12
 
Thanks and regards,
Kushagra